Producer

Suppose

Firm Production Function Q=50 à ƒ ƒ ¢ ˆ šLK

MPLE 25 VE MPK = 25à ƒ ƒ ¢ ˆ š2/

suppose

Price of Labor =W= P5.

Price of Capital = t = 120

Fand the cost minimizing input Combination if the firm wants to produce

1000 units.

Q=1600
1
Expert's answer
2021-10-20T16:42:54-0400

Cost minimizing combination of capital and labour is one where the marginal rate of substitution(MRTS) is equal to:


MRTS="\frac{MP_{L}}{MP_{K}}=\frac{w}{r}"


Marginal product of labour is "\frac{\delta Q}{\delta L}" =50 Ã ¢ K


Marginal product of capital= "\frac{\delta Q}{\delta K}" =50 Ã ¢ L


Therefore MRTS="\\frac{50 \u00c3\u0083 \u00a2 K}{50 \u00c3\u0083 \u00a2 L}=\\frac{K}{L}"


Set the marginal rate of technical substitution equal to the input price ratio to

determine the optimal capital-labor ratio.


"\frac{K}{L}=\frac{P5}{120} \therefore K=\frac{5PL}{120}" and "L=\frac{120K}{5P}"


Substitute for L in the production function and solve where K yields an output of 1000 units.


"1000=50\times\frac{120K}{5P}\times K"


"200=\frac{120K^{2}}{5P}"


"1000P=120K^{2}"


"K^{2}=8.33P \therefore K=(8.33P)^{0.5}"


To get L,we substitute K above in "L=\frac{120K}{5P}"


"L=\frac{120K}{5P}\times (8.33P)^{0.5}"


"L=\frac {24(8.33P)^{0.5}}{P}"


"\therefore K=(8.33P)^{0.5}" and "L=\frac{24(8.33P)^{0.5}}{P}" are the cost-minimizing levels of K and L.